OpenStudy (anonymous):

limit as x approaches e (from the right) of pi^(1/x-e)

OpenStudy (anonymous):

\[\lim_{x \rightarrow e^+} \pi^ (1/x-e)\]

OpenStudy (zarkon):

this \[\large\lim_{x\to e^{+}}\pi^{\frac{1}{x-e}}\] ?

OpenStudy (anonymous):

0